# StableToolBench: Towards Stable Large-Scale Benchmarking on Tool Learning of Large Language Models

### v2

@@ 6 Conclusion
In this paper, we propose StableToolBench, a benchmark developed to enhance the stability of ToolBench. Our analysis identified instability issues in evaluation processes of ToolBench and API status, causing variability in model performance assessments. To address this, we implement a caching system for consistent data availability. We also replace the real API server with an LLM simulated virtual server for reliable API behaviour simulation. Experiments show that StableToolBench significantly improves the stability of model performance evaluations, with the simulated APIs offering realism and the caching system contributing greatly to the enhanced stability of the benchmark.
@@ Appendix D Call Error Identification and Cache Filtering Rule
System
Imagine you are an API Server operating within a specialized tool, which contains a collection of distinct APIs. Your role is to deeply understand the function of each API based on their descriptions in the API documentation. As you receive specific inputs for individual API calls within this tool, analyze these inputs to determine their intended purpose. Your task is to craft a JSON formatted response that aligns with the expected output of the API, guided by the provided examples. Your responses must adhere to a specific JSON structure, which is as follows: { ‘‘error’’: ‘‘’’, ‘‘response’’: ‘‘Your_Response’’ } The error field should remain empty, indicating no errors in processing. The response field should contain the content you formulate based on the API’s functionality and the input provided. Ensure that your responses are meaningful, directly addressing the API’s intended functionality. If the provided examples are mostly error messages or lack substantial content, use your judgment to create relevant and accurate responses. The key is to maintain the JSON format’s integrity while ensuring that your response is an accurate reflection of the API’s intended output within the tool. Please note that your answer should not contain anything other than a json format object, which should be parsable directly to json. Note that: - your response should be around 100 to 200 words, containing rich information given the api input parameters. Keep Your answer short and simple. - your response must be effective and have practical content. - if the api response example if null or ineffective, ignore the example and give your independent response.
User
API Documentation: Documentation JSON file API Examples: Example input 1: Example response 1 Example input 2: Example response 2 Example input 3: Example response 3 API input: Argument JSON string, e.g: {‘‘category’’:‘‘Logistics’’,‘‘tool_name’’: ‘‘SQUAKE’’, ‘‘api_name’’: ‘‘Checkhealth’’,‘‘tool_input’’: ‘‘{}’’, ‘‘strip’’: ‘‘filter’’}
@@ Appendix H Prompt Used to Make API Calls
System
Imagine you are an API requester, Your role is to deeply understand the function of each API based on their descriptions in the API documentation. Your task is to craft a JSON formatted input that aligns with the expected input of the API, guided by the provided examples. Your responses must adhere to a specific JSON structure, which is as follows: Please note that your answer should not contain anything other than a json format object, which should be parsable directly to json. Note that: - your response should be around 100 to 500 words, containing rich information given the api input parameters. - your response must be effective and have practical content. - if the api response example if null or ineffective, ignore the example and give your independent response.
User
API Documentation: Documentation JSON file API Examples (if available): Example input 1: Example response 1 Example input 2: Example response 2 Example input 3: Example response 3 one more API Input example:

### v4

@@ 6 Conclusion
In this paper, we propose StableToolBench, a benchmark developed to enhance the stability of ToolBench. Our analysis identified instability issues in the evaluation processes of ToolBench and API status, causing variability in model performance assessments. To address this, we implement a caching system for consistent data availability. We also replace the real API server with an LLM-simulated virtual server for reliable API behaviour simulation. Experiments show that StableToolBench significantly improves the stability of model performance evaluations, with the simulated APIs offering realism and the caching system contributing greatly to the enhanced stability of the benchmark.
Acknowledgement
This work is supported by the National Natural Science Foundation of China (No. 62276152, 61925601). We also extend our gratitude to Jingwen Wu and Yao Li for their assistance with human evaluation and additional suggestions.
@@ Appendix D Call Error Identification and Cache Filtering Rule
System Imagine you are an API Server operating within a specialized tool, which contains a collection of distinct APIs. Your role is to deeply understand the function of each API based on their descriptions in the API documentation. As you receive specific inputs for individual API calls within this tool, analyze these inputs to determine their intended purpose. Your task is to craft a JSON formatted response that aligns with the expected output of the API, guided by the provided examples. Your responses must adhere to a specific JSON structure, which is as follows: { ‘‘error’’: ‘‘’’, ‘‘response’’: ‘‘Your_Response’’ } The error field should remain empty, indicating no errors in processing. The response field should contain the content you formulate based on the API’s functionality and the input provided. Ensure that your responses are meaningful, directly addressing the API’s intended functionality. If the provided examples are mostly error messages or lack substantial content, use your judgment to create relevant and accurate responses. The key is to maintain the JSON format’s integrity while ensuring that your response is an accurate reflection of the API’s intended output within the tool. Please note that your answer should not contain anything other than a json format object, which should be parsable directly to json. Note that: - your response should be around 100 to 200 words, containing rich information given the api input parameters. Keep Your answer short and simple. - your response must be effective and have practical content. - if the api response example if null or ineffective, ignore the example and give your independent response.
User API Documentation: Documentation JSON file API Examples: Example input 1: Example response 1 Example input 2: Example response 2 Example input 3: Example response 3 API input: Argument JSON string, e.g: {‘‘category’’:‘‘Logistics’’,‘‘tool_name’’: ‘‘SQUAKE’’, ‘‘api_name’’: ‘‘Checkhealth’’,‘‘tool_input’’: ‘‘{}’’, ‘‘strip’’: ‘‘filter’’}
@@ Appendix H Prompt Used to Make API Calls
System Imagine you are an API requester, Your role is to deeply understand the function of each API based on their descriptions in the API documentation. Your task is to craft a JSON formatted input that aligns with the expected input of the API, guided by the provided examples. Your responses must adhere to a specific JSON structure, which is as follows: Please note that your answer should not contain anything other than a json format object, which should be parsable directly to json. Note that: - your response should be around 100 to 500 words, containing rich information given the api input parameters. - your response must be effective and have practical content. - if the api response example if null or ineffective, ignore the example and give your independent response.
User API Documentation: Documentation JSON file API Examples (if available): Example input 1: Example response 1 Example input 2: Example response 2 Example input 3: Example response 3 one more API Input example: